# Dequantified Diffusion-Schrödinger Bridge for Density Ratio Estimation

### v2

@@ Appendix A Proofs
$=\int_{0}^{1}\int_{0}^{1}\mathbb{E}_{q_{s}^{\prime}}\left[\partial_{s}\log q_{s}^{\prime}\left[\mathbb{E}_{q_{t}^{\prime}}\left[\partial_{t}\log q_{t}^{\prime}\right]+e^{-\frac{|t-s|}{\gamma^{2}}}\left[\partial_{s}\log q_{s}^{\prime}-\mathbb{E}_{q_{s}^{\prime}}\left[\partial_{s}\log q_{s}^{\prime}\right]\right]+\mathcal{O}(\gamma^{4})\right]\right]\mathrm{d}t\mathrm{d}s\quad(\boldsymbol{\star})$
$=\int_{0}^{1}\int_{0}^{1}\left[\mathbb{E}_{q_{t}^{\prime}}\left[\partial_{t}\log q_{t}^{\prime}\right]\mathbb{E}_{q_{s}^{\prime}}\left[\partial_{s}\log q_{s}^{\prime}\right]+e^{-\frac{|t-s|}{\gamma^{2}}}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\right]\mathrm{d}t\mathrm{d}s+\mathcal{O}(\gamma^{4})$
$=\left(\int_{0}^{1}\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]\mathrm{d}t\right)\left(\int_{0}^{1}\mathbb{E}_{q_{s}^{\prime}}[\partial_{s}\log q_{s}^{\prime}]\mathrm{d}s\right)+\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\left(\int_{0}^{1}e^{-\frac{|t-s|}{\gamma^{2}}}\mathrm{d}s\right)\mathrm{d}t+\mathcal{O}(\gamma^{4})$
$=\left(\int_{0}^{1}\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]\mathrm{d}t\right)^{2}+\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\left(\gamma^{2}\left(2-e^{-\frac{t}{\gamma^{2}}}-e^{-\frac{1-t}{\gamma^{2}}}\right)\right)\mathrm{d}t+\mathcal{O}(\gamma^{4})$
$\leq\left(\int_{0}^{1}\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]\mathrm{d}t\right)^{2}+2\gamma^{2}\left(1-e^{-\frac{1}{2\gamma^{2}}}\right)\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\mathrm{d}t+\mathcal{O}(\gamma^{4})\quad(\boldsymbol{\star}\boldsymbol{\star}).$
$(\boldsymbol{\star})$ The conditional expectation $\mathbb{E}_{q_{t}^{\prime}}\left[\partial_{t}\log q_{t}^{\prime}\big{|}\mathbf{X}_{s}^{\prime}\right]$ satisfies the backward equation associated with the Markov semigroup. For small $\gamma$, the process resembles an Ornstein-Uhlenbeck process near equilibrium, where the covariance decays exponentially with the time separation $|t-s|$ due to the Chapman-Kolmogorov equation.
$(\boldsymbol{\star}\boldsymbol{\star})$ This inequality holds because the term $2-e^{-\frac{t}{\gamma^{2}}}-e^{-\frac{1-t}{\gamma^{2}}}$ is positive definite for $t\in[0,1]$ and attains its maxima at $t=0.5$ and minimum at $t=0$ and $t=1$. The first term is bounded by $C_{0}=\mathbb{E}_{\pi}\left[\left|\log\frac{q_{1}^{\prime}}{q_{0}^{\prime}}\right|^{2}\right]$ according to the Fubini’s theorem and the inequality $\mathbb{E}[\mathbf{X}]^{2}\leq\mathbb{E}[\mathbf{X}^{2}]$.
@@
$\mathbb{E}_{\pi}[\mathcal{E}^{2}]\leq 4\gamma^{2}\left(1-e^{-\frac{1}{2\gamma^{2}}}\right)\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\mathrm{d}t+4C_{0}+\mathcal{O}(\gamma^{4}).$ (44)
@@
$\mathbb{E}_{\pi}[\mathcal{E}^{2}]$ $\leq\mathcal{V}=4\gamma^{2}\left(1-e^{-\frac{1}{2\gamma^{2}}}\right)\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\mathrm{d}t+4C_{0}+\mathcal{O}(\gamma^{4})$ (49)
$=4\gamma^{2}\left(1-e^{-\frac{1}{2\gamma^{2}}}\right)\int_{0}^{1}\left[\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]+\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}\right]\mathrm{d}t+4C_{0}+\mathcal{O}(\gamma^{4})$
$=4\gamma^{2}\left(1-e^{-\frac{1}{2\gamma^{2}}}\right)\left[\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]\int_{0}^{1}\frac{1}{\sigma_{t}^{2}}\mathrm{d}t+\int_{0}^{1}\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}\mathrm{d}t\right]+4C_{0}+\mathcal{O}(\gamma^{4})$
Thus, the difference between the upper bound of $\mathbb{E}_{\pi}[\mathcal{E}^{2}]$ for DSBI and DDBI becomes:
$\mathcal{V}_{\text{DDBI}}-\mathcal{V}_{\text{DSBI}}$ $=4\gamma^{2}\underbrace{\left(1-e^{-\frac{1}{2\gamma^{2}}}\right)}_{>0}\underbrace{\left[\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]-\mathbb{E}_{\pi_{2\gamma^{2}}}\left[\left\|\hat{\mathbf{X}}_{0}^{\prime}-\hat{\mathbf{X}}_{1}^{\prime}\right\|^{2}\right]\right]}_{\geq 0(\boldsymbol{\star}\boldsymbol{\star}\boldsymbol{\star})}\int_{0}^{1}\frac{1}{\sigma_{t}^{2}}\mathrm{d}t+\mathcal{O}(\gamma^{4}).$ (50)
@@ B.1 Special Cases of DI
In the case of TRE, the interpolation strategy, as detailed in Eq. (2), represents a specific case of $\mathbf{I}$, characterized by $\alpha_{t}=\sqrt{1-\eta_{t}^{2}},\beta_{t}=\eta_{t}$, with $t$ taking discrete values $10,1/M,2/M,\ldots,$ 1. For DRE-$\infty$, the coefficients are defined as $\alpha_{t}=\exp\{-0.25(\beta_{\max}-\beta_{\min})t^{2}-0.5\beta_{\min}t\}$ and $\beta_{t}=\sqrt{1-\alpha_{t}^{2}}$ for the MNIST dataset and $\alpha_{t}=1-t$ and $\beta_{t}=t$ for other datasets. The corresponding stochastic process for the former one aligns with the solution to variance preserving (VP) SDEs (Song et al., 2020b; Li et al., 2024b; Xin et al., 2024; Zhou et al., 2025b).
@@ B.2 From Optimal Transport to Entropic Regularization
where $\mathbf{X}_{0}\sim q_{0}$ and $\mathbf{X}_{1}\sim q_{1}$. This interpolation aligns with the Benamou-Brenier formulation (McCann, 1997; Zhou et al., 2024), where the transport paths minimize the kinetic energy in the space of probability measures.
The natural connections between optimal transport theory and straight-line interpolations motivate the concept of Batch Optimal Transport (BatchOT) (Pooladian et al., 2023; zhao2025less). BatchOT provides a pseudo-deterministic coupling mechanism by extending the OT principles to minibatch sampling. This ensures practical scalability and aligns theoretical transport paths with computational requirements.
@@ C.1 Comparison of the Trajectory Sets for Interpolation Strategies
(b) DBI: DBI introduces stochasticity through Brownian Bridge noise, expanding support and enabling broader trajectory exploration (Figure 1(b)). Compared to DI, DBI provides greater coverage and variability while retaining tractability, reducing the rigidity of interpolation paths (zhao2023learning).
@@ C.4 Density Estimation
Energy-based Modeling on MNIST. We applied the proposed $\text{D}^{3}\text{RE}$ framework to density estimation on the MNIST (LeCun, 1998) dataset, leveraging pre-trained energy-based models (EBMs) (Choi et al., 2022). Let $q_{1}(\mathbf{x})$ denote the MNIST data distribution and $q_{0}(\mathbf{x})$ a simple noise distribution with three different settings, as reported in (Rhodes et al., 2020): Gaussian noise, Gaussian copula, and Rational Quadratic Neural Spline Flow (RQ-NSF) (Durkan et al., 2019). We applied an modified version of DDBI of the form $\mathbf{X}_{t}^{\prime}=\alpha_{t}\mathbf{X}_{0}+\beta_{t}\text{EBM}(\mathbf{X}_{1})+\sqrt{t(1-t)\gamma^{2}+(\alpha_{t}^{2}+\beta_{t}^{2})\varepsilon}\mathbf{Z}_{t}$, where $\mathbf{X}_{0}\sim q_{0}(\mathbf{x}),\mathbf{X}_{1}\sim q_{1}(\mathbf{x})$, $\mathbf{Z}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{E}_{d})$, $\alpha_{t}=\exp\{-0.25(\beta_{\max}-\beta_{\min})t^{2}-0.5\beta_{\min}t\}$ and $\beta_{t}=\sqrt{1-\alpha_{t}^{2}}$. $\beta_{\min}$ and $\beta_{\max}$ are set to 0.1 and 20, respectively. The results are reported in bits-per-dimension (BPD), evaluated as $\text{BPD}=-\frac{1}{d\ln{2}}\mathbb{E}_{\mathbf{x}\sim q_{1}(\mathbf{x})}\left[\log q_{1}(\mathbf{x})\right]$, where the expectation reflects the log-density of the MNIST dataset. Exact BPD computation is infeasible for EBMs; therefore, we estimate it using two annealed MCMC methods: Annealed Importance Sampling (AIS) (Neal, 2001) and Reverse Annealed Importance Sampling Estimator (RAISE) (Burda et al., 2015).
@@ C.6 Ablation Study on OTR
To evaluate the role of OTR in the model’s performance, we conducted an ablation study comparing models without OTR (baseline, DI), models with OTR (DI+OTR), and models from the $\text{D}^{3}\text{RE}$ framework (DDBI and DSBI).
In Figure 2, DI generates distorted and misaligned intermediate distributions. This shows its limited ability to align with the target distribution. DI+OTR improves alignment significantly and produces more accurate patterns. However, some details remain suboptimal. Models from the $\text{D}^{3}\text{RE}$ framework further enhance distribution quality. Among them, DSBI achieves the most precise alignment with the target distribution, both globally and locally. These results highlight the critical role of OTR in improving the alignment of intermediate distributions.
@@
Number of Function Evaluations. We analyze the impact of OTR on NFE, noting that DI and DDBI do not utilize OTR. Our observations show that applying OTR significantly reduces NFE. Figure 6 compares NFE across four methods in DRE, highlighting substantial variations in computational efficiency. The first approach exhibits the highest NFE, indicating reliance on iterative procedures requiring repeated function evaluations. The second approach achieves a moderate reduction in NFE, likely by minimizing redundant evaluations through minimized transport costs. More results can be found in Table 4.
Table 4: Comparison of the NFE for different density ratio settings. The methods with OTR have lower NFE to those without OTR.
Source Target DI DI+OTR (ours) DDBI (ours) DSBI (ours)
$\mathsf{pinwheel}$ $\mathsf{2spirals}$ 572 451 609 500
$\mathsf{8gaussians}$ $\mathsf{2spirals}$ 506 415 508 454
$\mathsf{8gaussians}$ $\mathsf{pinwheel}$ 521 397 524 446
$\mathsf{moons}$ $\mathsf{2spirals}$ 526 463 530 476
$\mathsf{moons}$ $\mathsf{8gaussians}$ 475 410 444 422
$\mathsf{moons}$ $\mathsf{pinwheel}$ 541 445 545 467
$\mathsf{circles}$ $\mathsf{2spirals}$ 532 456 518 438
$\mathsf{circles}$ $\mathsf{8gaussians}$ 481 402 432 384
$\mathsf{circles}$ $\mathsf{moons}$ 501 450 454 405
$\mathsf{circles}$ $\mathsf{pinwheel}$ 548 438 534 429
$\mathsf{swissroll}$ $\mathsf{2spirals}$ 535 422 539 482
$\mathsf{swissroll}$ $\mathsf{8gaussians}$ 484 369 453 428
$\mathsf{swissroll}$ $\mathsf{circles}$ 510 409 463 411
$\mathsf{swissroll}$ $\mathsf{moons}$ 504 417 475 449
$\mathsf{swissroll}$ $\mathsf{pinwheel}$ 550 404 554 473

### v3

@@ Appendix A Proofs
$=\int_{0}^{1}\int_{0}^{1}\mathbb{E}_{q_{s}^{\prime}}\left[\partial_{s}\log q_{s}^{\prime}\left[\mathbb{E}_{q_{t}^{\prime}}\left[\partial_{t}\log q_{t}^{\prime}\right]+e^{-\frac{|t-s|}{\tau^{2}}}\left[\partial_{s}\log q_{s}^{\prime}-\mathbb{E}_{q_{s}^{\prime}}\left[\partial_{s}\log q_{s}^{\prime}\right]\right]+\mathcal{O}(\tau^{4})\right]\right]\mathrm{d}t\mathrm{d}s\quad(\boldsymbol{\star})$
$=\int_{0}^{1}\int_{0}^{1}\left[\mathbb{E}_{q_{t}^{\prime}}\left[\partial_{t}\log q_{t}^{\prime}\right]\mathbb{E}_{q_{s}^{\prime}}\left[\partial_{s}\log q_{s}^{\prime}\right]+e^{-\frac{|t-s|}{\tau^{2}}}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\right]\mathrm{d}t\mathrm{d}s+\mathcal{O}(\tau^{4})$
$=\left(\int_{0}^{1}\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]\mathrm{d}t\right)\left(\int_{0}^{1}\mathbb{E}_{q_{s}^{\prime}}[\partial_{s}\log q_{s}^{\prime}]\mathrm{d}s\right)+\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\left(\int_{0}^{1}e^{-\frac{|t-s|}{\tau^{2}}}\mathrm{d}s\right)\mathrm{d}t+\mathcal{O}(\tau^{4})$
$=\left(\int_{0}^{1}\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]\mathrm{d}t\right)^{2}+\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\left(\tau^{2}\left(2-e^{-\frac{t}{\tau^{2}}}-e^{-\frac{1-t}{\tau^{2}}}\right)\right)\mathrm{d}t+\mathcal{O}(\tau^{4})$
$\leq\left(\int_{0}^{1}\mathbb{E}_{q_{t}^{\prime}}[\partial_{t}\log q_{t}^{\prime}]\mathrm{d}t\right)^{2}+2\tau^{2}\left(1-e^{-\frac{1}{2\tau^{2}}}\right)\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\mathrm{d}t+\mathcal{O}(\tau^{4})\quad(\boldsymbol{\star}\boldsymbol{\star}).$
$(\boldsymbol{\star})$ The conditional expectation $\mathbb{E}_{q_{t}^{\prime}}\left[\partial_{t}\log q_{t}^{\prime}\big{|}\mathbf{X}_{s}^{\prime}\right]$ satisfies the backward equation associated with the Markov semigroup. For small $\tau$, the process resembles an Ornstein-Uhlenbeck process near equilibrium, where the covariance decays exponentially with the time separation $|t-s|$ due to the Chapman-Kolmogorov equation.
$(\boldsymbol{\star}\boldsymbol{\star})$ This inequality holds because the term $2-e^{-\frac{t}{\tau^{2}}}-e^{-\frac{1-t}{\tau^{2}}}$ is strictly positive for $t\in[0,1]$ and attains its maxima at $t=0.5$ and minimum at $t=0$ and $t=1$. The first term is bounded by $C_{0}=\mathbb{E}_{\pi}\left[\left|\log\frac{q_{1}^{\prime}}{q_{0}^{\prime}}\right|^{2}\right]$ according to the Fubini’s theorem and the inequality $\mathbb{E}[\mathbf{X}]^{2}\leq\mathbb{E}[\mathbf{X}^{2}]$.
@@
$\mathbb{E}_{\pi}[\mathcal{E}^{2}]\leq 4\tau^{2}\left(1-e^{-\frac{1}{2\tau^{2}}}\right)\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\mathrm{d}t+4C_{0}+\mathcal{O}(\tau^{4}).$ (44)
@@
$\mathbb{E}_{\pi}[\mathcal{E}^{2}]$ $\leq 2\mathcal{V}+2C_{0}=4\tau^{2}\left(1-e^{-\frac{1}{2\tau^{2}}}\right)\int_{0}^{1}\text{Var}_{q_{t}^{\prime}}(\partial_{t}\log q_{t}^{\prime})\mathrm{d}t+4C_{0}+\mathcal{O}(\tau^{4})$ (49)
$=4\tau^{2}\left(1-e^{-\frac{1}{2\tau^{2}}}\right)\int_{0}^{1}\left[\frac{1}{\sigma_{t}^{2}}\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]+\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}\right]\mathrm{d}t+4C_{0}+\mathcal{O}(\tau^{4})$
$=4\tau^{2}\left(1-e^{-\frac{1}{2\tau^{2}}}\right)\left[\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]\int_{0}^{1}\frac{1}{\sigma_{t}^{2}}\mathrm{d}t+\int_{0}^{1}\frac{\dot{\sigma}_{t}^{4}d}{2\sigma_{t}^{4}}\mathrm{d}t\right]+4C_{0}+\mathcal{O}(\tau^{4}),$
when $\alpha_{t}=1-t$ and $\beta_{t}=t$. Thus, the difference between the upper bound of $\mathbb{E}_{\pi}[\mathcal{E}^{2}]$ for DSBI and DDBI becomes:
$\mathcal{V}_{\text{DDBI}}-\mathcal{V}_{\text{DSBI}}$ $=4\tau^{2}\underbrace{\left(1-e^{-\frac{1}{2\tau^{2}}}\right)}_{>0}\underbrace{\left[\mathbb{E}_{\pi}\left[\left\|\mathbf{X}_{0}^{\prime}-\mathbf{X}_{1}^{\prime}\right\|^{2}\right]-\mathbb{E}_{\pi_{2\gamma^{2}}}\left[\left\|\hat{\mathbf{X}}_{0}^{\prime}-\hat{\mathbf{X}}_{1}^{\prime}\right\|^{2}\right]\right]}_{\geq 0(\boldsymbol{\star}\boldsymbol{\star}\boldsymbol{\star})}\int_{0}^{1}\frac{1}{\sigma_{t}^{2}}\mathrm{d}t+\mathcal{O}(\tau^{4}).$ (50)
@@ B.1 Special Cases of DI
In the case of TRE, the interpolation strategy, as detailed in Eq. (2), represents a specific case of $\mathbf{I}$, characterized by $\alpha_{t}=\sqrt{1-\eta_{t}^{2}},\beta_{t}=\eta_{t}$, with $t$ taking discrete values $10,1/M,2/M,\ldots,$ 1. For DRE-$\infty$, the coefficients are defined as $\alpha_{t}=\exp\{-0.25(\beta_{\max}-\beta_{\min})t^{2}-0.5\beta_{\min}t\}$ and $\beta_{t}=\sqrt{1-\alpha_{t}^{2}}$ for the MNIST dataset and $\alpha_{t}=1-t$ and $\beta_{t}=t$ for other datasets. The corresponding stochastic process for the former one aligns with the solution to variance preserving (VP) SDEs (Song et al., 2021; Li et al., 2024b; Xin et al., 2024; Zhou et al., 2025b).
@@ B.2 From Optimal Transport to Entropic Regularization
where $\mathbf{X}_{0}\sim q_{0}$ and $\mathbf{X}_{1}\sim q_{1}$. This interpolation aligns with the Benamou-Brenier formulation (McCann, 1997; Zhou et al., 2024b), where the transport paths minimize the kinetic energy in the space of probability measures.
The natural connections between optimal transport theory and straight-line interpolations motivate the concept of Batch Optimal Transport (BatchOT) (Pooladian et al., 2023; Zhao et al., 2025). BatchOT provides a pseudo-deterministic coupling mechanism by extending the OT principles to minibatch sampling. This ensures practical scalability and aligns theoretical transport paths with computational requirements.
@@ C.1 Comparison of the Trajectory Sets for Interpolation Strategies
(b) DBI: DBI introduces stochasticity through Brownian Bridge noise, expanding support and enabling broader trajectory exploration (Figure 1(b)). Compared to DI, DBI provides greater coverage and variability while retaining tractability, reducing the rigidity of interpolation paths (Zhao et al., 2023).
@@ C.4 Density Estimation
Energy-based Modeling on MNIST. We applied the proposed $\text{D}^{3}\text{RE}$ framework to density estimation on the MNIST dataset, leveraging pre-trained energy-based models (EBMs) (Choi et al., 2022). Let $q_{1}(\mathbf{x})$ denote the MNIST data distribution and $q_{0}(\mathbf{x})$ a simple noise distribution with three different settings, as reported in (Rhodes et al., 2020): Gaussian noise, Gaussian copula, and Rational Quadratic Neural Spline Flow (RQ-NSF) (Durkan et al., 2019). We applied an modified version of DDBI of the form $\mathbf{X}_{t}^{\prime}=\alpha_{t}\mathbf{X}_{0}+\beta_{t}\text{EBM}(\mathbf{X}_{1})+\sqrt{t(1-t)\gamma^{2}+(\alpha_{t}^{2}+\beta_{t}^{2})\varepsilon}\mathbf{Z}_{t}$, where $\mathbf{X}_{0}\sim q_{0}(\mathbf{x}),\mathbf{X}_{1}\sim q_{1}(\mathbf{x})$, $\mathbf{Z}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{E}_{d})$, $\alpha_{t}=\exp\{-0.25(\beta_{\max}-\beta_{\min})t^{2}-0.5\beta_{\min}t\}$ and $\beta_{t}=\sqrt{1-\alpha_{t}^{2}}$. $\beta_{\min}$ and $\beta_{\max}$ are set to 0.1 and 20, respectively. The results are reported in bits-per-dimension (BPD), evaluated as $\text{BPD}=-\frac{1}{d\ln{2}}\mathbb{E}_{\mathbf{x}\sim q_{1}(\mathbf{x})}\left[\log q_{1}(\mathbf{x})\right]$, where the expectation reflects the log-density of the MNIST dataset. Exact BPD computation is infeasible for EBMs; therefore, we estimate it using two annealed MCMC methods: Annealed Importance Sampling (AIS) (Neal, 2001) and Reverse Annealed Importance Sampling Estimator (RAISE) (Burda et al., 2015).
@@ C.6 Ablation Study on OTR
Number of Function Evaluations. We analyze the impact of OTR on NFE, noting that DI and DDBI do not utilize OTR. Our observations show that applying OTR significantly reduces NFE. Figure 8 compares NFE across four methods in DRE, highlighting substantial variations in computational efficiency. The first approach exhibits the highest NFE, indicating reliance on iterative procedures requiring repeated function evaluations. The second approach achieves a moderate reduction in NFE, likely by minimizing redundant evaluations through minimized transport costs.